OpenStudy (anonymous):

Use the quadratic formula to find the real solutions, if any, to the equation. x^2 + 2x - 5 = 0

OpenStudy (anonymous):

I'm stuck at \[=\frac{ -2 \pm \sqrt{-5} }{ 2 }\]

OpenStudy (helder_edwin):

for the equation \(ax^2+bx+c=0\) u have solutions \[ \large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \]

OpenStudy (anonymous):

Mhmm o;

OpenStudy (helder_edwin):

\[ \large b ^2-4ac=4-4(1)(-5)= \]

OpenStudy (anonymous):

Look at my first comment xD

OpenStudy (helder_edwin):

i did and i'm correcting u.

OpenStudy (anonymous):

It's -5

OpenStudy (anonymous):

OH. it's 0.

OpenStudy (anonymous):

so that means no solutions?

OpenStudy (helder_edwin):

\[ \large 4-4(1)(-5)=4+20=24 \]

OpenStudy (anonymous):

Oh, yeah. :c that was nooby sorry.

OpenStudy (helder_edwin):

so \[ \large x=\frac{-2\pm\sqrt{24}}{2} \]

OpenStudy (anonymous):

What would you do about the -2 and 2?

OpenStudy (helder_edwin):

\[ \large \sqrt{24}=\sqrt{6\cdot4}=2\sqrt{6} \]

OpenStudy (helder_edwin):

\[ \large x=\frac{-2\pm2\sqrt{6}}{2}=-1\pm\sqrt{6} \]

OpenStudy (anonymous):

Ohhh. So would the solutions be \[-1+\sqrt{6}, -1-\sqrt{6}\]

OpenStudy (helder_edwin):

yes

OpenStudy (anonymous):

Thank you so much!

OpenStudy (helder_edwin):

u r welcome